OpenStudy (anonymous):

Determine which points lie on the line that contains the point (2,3) and has a slope of -7/4

OpenStudy (anonymous):

Hint: Put into point-slope form: y - y1 = m(x - x1) m = slope y1 and x1 = values of x and y y - 3 = -7/4(x - 2) Solve for y in order to convert to slope-intercept form. y = mx + b where m = slope b = y-intercept. To determine which of the following points also lies on the line, simply substitute the value of x and see if the resulting y is the same as the y of the point in question.

OpenStudy (anonymous):

@andrew.gordon